lisa has collected data to find that the number of pages per book on a book shelf has a normal dostributuon. what is fhe probability that a randomly selected book has fewer than 167 pages if the mean is196 pages and fhe standard deviation is 28 pages? use the empirical rule. enter your answer as a percent rounded to two decimal placed if necessay.
Answer #1

z score for 167 = (167 - 196)/28 = -1.04

Hence,

P(Fewer than 167 pages)

= P(z < -1.04)

= 0.1502

= 15.02 %